# VisionTS: Visual Masked Autoencoders Are Free-Lunch Zero-Shot Time Series Forecasters

### v3

@@ Abstract
Foundation models have emerged as a promising approach in time series forecasting (TSF). Existing approaches either repurpose large language models (LLMs) or build large-scale time series datasets to develop TSF foundation models for universal forecasting. However, these methods face challenges due to the severe cross-domain gap or in-domain heterogeneity. This paper explores a new road to building a TSF foundation model from rich, high-quality natural images. Our key insight is that a visual masked autoencoder, pre-trained on the ImageNet dataset, can naturally be a numeric series forecaster. By reformulating TSF as an image reconstruction task, we bridge the gap between image pre-training and TSF downstream tasks. Surprisingly, without further adaptation in the time series domain, the proposed VisionTS could achieve better zero-shot forecast performance than existing TSF foundation models. With fine-tuning for one epoch, VisionTS could further improve the forecasting and achieve state-of-the-art performance in most cases. Extensive experiments reveal intrinsic similarities between images and real-world time series, suggesting that visual models may offer a “free lunch” for TSF and highlight the potential for future cross-modality research. Our code is available in the Supplementary Material.
@@
Figure 1: Long-term forecasting performance comparison. Our VisionTS, without any training on time series data, outperforms the pure time series foundation model MoiraiLarge in the zero-shot setting.
@@ 3 Methodology
Given a univariate input $X\in\mathbb{R}^{L}$, the first goal is to transform it into a 2D matrix. We propose to segment it into $\lfloor\nicefrac{{L}}{{P}}\rfloor$ subsequences of length $P$, where $P$ is the periodicity. Notably, when the time series lacks clear periodicity, we can set $P=1$ directly, which is also effective in our experiments (Section C.5). In practice, $P$ can be determined using statistical methods like Fast Fourier Transform (Wu et al., 2023; Chen et al., 2024) or domain knowledge like sampling frequency (Godahewa et al., 2021; Alexandrov et al., 2020). In this paper, we select $P$ based on the sampling frequency, elaborated in Section B.2.
@@
MAE standardizes each image based on the mean and standard deviation computed on ImageNet. Therefore, we apply instance normalization to ${\bm{I}}_{\text{raw}}$, which is also a standard practice in current TSF (Kim et al., 2022). Notably, we observed that normalizing ${\bm{I}}_{\text{raw}}$ to a standard deviation of $r$, where $r$ is a hyperparameter less than 1, yields superior performance. One explanation is that the magnitude of inputs/outputs during MAE pretraining is constrained by the limited range of color values. Therefore, reducing the magnitude of ${\bm{I}}_{\text{raw}}$ prevents exceeding these limits. However, an excessively low $r$ can result in values that are difficult to distinguish. We found that a moderate value (0.4) of $r$ performs well across most scenarios (See Section C.8 for more details). Let ${\bm{I}}_{\text{norm}}$ denote the normalized matrix, which is computed as follows:
@@
Moreover, we found that reducing the width of the visible portion can further improve performance. One possible explanation is that MAE uses a large masked ratio during pre-training, with only 25% of patches visible. Reducing the image width may align the masked ratio more closely with pre-training. Therefore, we propose multiplying $n$ by a hyperparameter $c\in[0,1]$. Similar to $r$, we found that setting $c=0.4$ performs well in our experiments (See Section C.8). This can be formulated as $n=\left\lfloor c\cdot N\cdot\nicefrac{{L}}{{(L+H)}}\right\rfloor$.
@@
Table 1: Zero-shot or few-shot results on the long-term TSF benchmark. Results are averaged across prediction lengths {96, 192, 336, 720}, with full results in Table 9 (Section C.2). Bold: the best result.
@@ 4 Experiments
Figure 4: Performance on the GIFT-Eval Leaderboard.
Figure 5: Aggregated results on the Monash TSF Benchmark, with full results in Table 13 (Section C.5).
We follow the standard evaluation protocol proposed by Woo et al. (2024) to test our VisionTS on 35 widely-used TSF benchmarks, and additionally evaluate it on the GIFT-Eval (Aksu et al., 2024) which is the largest TSF benchmark for zero-shot foundation models. We use MAE (Base) as our backbone by default. Baseline and benchmark details are elaborated in Section B.1.
@@ 4.1 Zero-Shot Time Series Forecasting
We first evaluate VisionTS’s zero-shot TSF performance without fine-tuning on time-series modalities. To prevent data leakage, we selected six widely-used datasets from the long-term TSF benchmark that are not included in Moirai’s pre-training set for evaluation. Since most baselines cannot perform zero-shot forecasting, we report their few-shot results by fine-tuning on the 10% of the individual target datasets. We also evaluate the Monash benchmark and GIFT-Eval benchmark. Notably, the Monash benchmark is more challenging for VisionTS since they were used in Moirai’s pre-training but not for VisionTS. We set the hyperparameters to $r=c=0.4$. Following common practice (Zhou et al., 2023; Woo et al., 2024), we conduct hyperparameter tuning on validation sets to determine the optimal context length $L$, detailed in Section C.1.
Table 2: Average MSE of different MAE variants, with full results in Table 15 (Section C.6).
@@
Table 1 shows that VisionTS surprisingly achieves the best forecasting performance in most cases (7 out of 14). Specifically, VisionTS demonstrates a relative average MSE reduction of approximately 6% compared to MoiraiSmall and MoiraiLarge, and performs comparably to MoiraiBase. When compared to the various few-shot baselines, VisionTS shows a relative average MSE reduction ranging from 8% to 84%. Given that all baselines except for VisionTS are trained on the time-series domain, this result is particularly encouraging. It suggests that the transferability from images to time-series is stronger than from text to time-series, and even comparable to the in-domain transferability between time-series. We also include a comparison with two TSF foundation models, TimesFM (Das et al., 2024) and LLMTime (Gruver et al., 2023), in Section C.3, as well as traditional algorithms (ETS, ARIMA, and Seasonal Naïve) in Section C.4. Results show that VisionTS still outperforms all of these baselines.
Results on GIFT-Eval and Monash Benchmarks
Fig. 4 shows the comparison of VisionTS with six previously published TSF foundation models on the GIFT-Eval TSF Leaderboard [^2] https://huggingface.co/spaces/Salesforce/GIFT-Eval, where VisionTS surprisingly ranked first in terms of normalized MASE. Besides, Fig. 5 shows the results aggregated from 29 Monash datasets, showing that VisionTS in the zero-shot setting surpasses all models individually trained on each dataset and significantly outperforms the other cross-domain baseline (i.e., LLMTime). It achieves second place among all baselines, just behind Moirai that pre-trained on all the training datasets. This promising result highlights VisionTS’s strong zero-shot forecasting ability and effective cross-modality transferability.
@@ 4.2 Further Analysis of VisionTS
In Table 3 (full results in Section C.6), we observe that the overall performance of three MAE variants (112M, 330M, and 657M) outperforms MoiraiSmall and MoiraiLarge. Particularly, larger models show a slight decrease in performance. This may be due to larger visual models overfitting image-specific features, reducing their transferability. A similar phenomenon was reported in Moirai, where larger models were found to degrade performance. We leave the exploration of scaling laws in image-based TSF foundation models for the future. Additionally, to explore the potential with other vision models, we also test LaMa (Suvorov et al., 2022), a visual inpainting model. Results in Section C.6 demonstrate that VisionTS with LaMa performs similarly to Moirai in the zero-shot setting. This suggests that the performance is driven by the inherent similarity between images and time series, not solely by the MAE model.
@@
Table 4: Aggregated full-shot forecasting performance on eight long-term TSF benchmarks (ETTh1, ETTh2, ETTm1, ETTm2, Illness, Weather, Traffic, and Electricity). VisionTS is fine-tuned only a single epoch on each dataset except for Illness. Due to the space limit, we report the $1^{\text{st}}$ count for each baseline, with full results in Table 19 (Section D.2).
@@
Section C.8 illustrates the impact of three hyperparameters. For context length $L$, as shown in Fig. 7, performance typically improves with increasing $L$, particularly on high-frequency datasets like Weather (10-minute frequency) and ETTm1/ETTm2 (15-minute frequency). This aligns with other TSF foundation models like Moirai. As for the normalization constant $r$ and alignment constant $c$, when both of them are around 0.4, performance is generally well across most benchmarks.
@@
We conduct experiments to validate our choices in the Alignment step, detailed in Section C.7. First, we test three different interpolation strategies, which shows that Bilinear interpolation performs best. Second, we apply horizontal and vertical flips on the image to examine whether the assumed left-to-right, top-to-bottom order is efficient. Results show that these changes do not significantly affect performance, suggesting that image reconstruction is isotropic and not influenced by certain orientation.
@@
In Appendix E, we visualize the zero-shot forecasting of VisionTS alongside the input and reconstruction images, highlighting both successful cases (where VisionTS outperforms Moirai) and failures (where Moirai prevails). When the input exhibits strong regularity (Fig. 11), VisionTS effectively forecasts both the periodicity (via segmentation) and trends (via MAE’s capabilities). In contrast, Moirai, akin to seasonal naïve methods, struggles to capture inter-period trends. For less-structured input (Figs. 12, 13 and 14), Moirai adopts a conservative approach with lower volatility to minimize errors, while VisionTS takes a more aggressive stance. This strategy occasionally yields more accurate trend predictions (Figs. 12 and 13) but may also result in greater MAE (Fig. 14).
@@ 4.3 Full-Shot Long-Term Time Series Forecasting
We evaluate the full-shot capability of each baseline trained on individual long-term TSF benchmarks. In addition to the six datasets used for zero-shot forecasting, we also include the popular Traffic and Illness datasets. As self-attention and feed-forward layers contain rich knowledge that can be transferred to TSF, we choose to fine-tune only the layer normalization (LN) layers while freezing the other parameters, which is also adopted by Zhou et al. (2023). Training details are elaborated in Section D.1.
@@
Table 4 summarizes the full-shot results, with the full results and standard deviations detailed in Section D.2. It shows that VisionTS outperforms other baselines in most cases (46 out of 80), surpassing the non-pretrained PatchTST and the language-pretrained GPT4TS. Remarkably, except for Illness with the least data, VisionTS demands only a single epoch of fine-tuning. This suggests that even minimal fine-tuning enables VisionTS to adapt to time series effectively. Compared with Table 1, fine-tuning provides limited benefits for ETTh1 and ETTh2 but significantly improves other datasets. We attribute this to the smaller data scale of ETTh1 and ETTh2.
@@
Tan et al. (2024) proposed several ablation variants for text-based foundation models, including w/o LLM (removing the LLM), LLM2Attn/LLM2Trsf (replacing the LLM with a single self-attention/Transformer layer), and RandLLM (randomly initializing the LLM). They found no significant performance differences and concluded that textual knowledge is unnecessary for TSF. We conducted similar ablations to assess the role of the vision model (VM), including w/o VM, VM2Attn, VM2Trsf, and RandVM. Section D.3 shows that these variants lead to worse performance, indicating that visual knowledge is beneficial for TSF.
@@
As stated before, we fine-tune only the layer normalization (LN). We also tested fine-tuning the bias, MLP, or attention layers, in addition to full fine-tuning and freezing. All hyperparameters were kept constant. Note that freezing differs from the previous zero-shot experiment, where a longer context length was used. Section D.3 show that fine-tuning LN is the best. Modifying MLP or attention layers results in significant performance drops, suggesting that valuable knowledge resides in these components.
@@ 5 Related Work
Large Language Models (LLMs) pre-trained on large amounts of text data are being applied to TSF tasks. We review this line of work in Appendix A.
@@
Self-supervised pre-training a TSF model on the same dataset used for downstream TSF tasks is a well-explored topic (Ma et al., 2023; Zhang et al., 2024), such as denoising autoencoders (Zerveas et al., 2021) or contrastive learning (Woo et al., 2022a; Yue et al., 2022). They follow a similar paradigm to the masked autoencoder (MAE) in computer vision. However, these methods rarely examine the cross-dataset generalization capabilities. Recently, research has shifted towards training universal foundation models, by collecting large-scale time series datasets from diverse domains (Goswami et al., 2024; Liu et al., 2024; Das et al., 2024; Dong et al., 2024; Feng et al., 2024) or generating numerous synthetic time series data (Fu et al., 2024; Yang et al., 2024). As a representative method, Woo et al. (2024) collected 27 billion observations across nine domains and trained TSF foundation models of various scales, achieving strong zero-shot performance. However, given the severe heterogeneity, constructing high-quality large datasets poses significant challenges for building these foundation models.

### v4

@@ Abstract
Foundation models have emerged as a promising approach in time series forecasting (TSF). Existing approaches either repurpose large language models (LLMs) or build large-scale time series datasets to develop TSF foundation models for universal forecasting. However, these methods face challenges due to the severe cross-domain gap or in-domain heterogeneity. This paper explores a new road to building a TSF foundation model from rich, high-quality natural images. Our key insight is that a visual masked autoencoder, pre-trained on the ImageNet dataset, can naturally be a numeric series forecaster. By reformulating TSF as an image reconstruction task, we bridge the gap between image pre-training and TSF downstream tasks. Surprisingly, without further adaptation in the time series domain, the proposed VisionTS could achieve better zero-shot forecast performance than existing TSF foundation models. With fine-tuning for one epoch, VisionTS could further improve the forecasting and achieve state-of-the-art performance in most cases. Extensive experiments reveal intrinsic similarities between images and real-world time series, suggesting that visual models may offer a “free lunch” for TSF and highlight the potential for future cross-modality research. Our code is publicly available at https://github.com/Keytoyze/VisionTS.
@@
Figure 1: Long-term forecasting (left) and GIFT-Eval (right) performance comparison. Our VisionTS, without any training on time series data, outperforms the pure time series foundation models in the zero-shot setting.
@@ 3 Methodology
Given a univariate input $X\in\mathbb{R}^{L}$, the first goal is to transform it into a 2D matrix. We propose to segment it into $\lfloor\nicefrac{{L}}{{P}}\rfloor$ subsequences of length $P$, where $P$ is the periodicity. Notably, when the time series lacks clear periodicity, we can set $P=1$ directly, which is also effective in our experiments (Section B.6). In practice, $P$ can be determined using statistical methods like Fast Fourier Transform (Wu et al., 2023; Chen et al., 2024) or domain knowledge like sampling frequency (Godahewa et al., 2021; Alexandrov et al., 2020). In this paper, we select $P$ based on the sampling frequency, elaborated in Section A.2.
@@
MAE standardizes each image based on the mean and standard deviation computed on ImageNet. Therefore, we apply instance normalization to ${\bm{I}}_{\text{raw}}$, which is also a standard practice in current TSF (Kim et al., 2022). Notably, we observed that normalizing ${\bm{I}}_{\text{raw}}$ to a standard deviation of $r$, where $r$ is a hyperparameter less than 1, yields superior performance. One explanation is that the magnitude of inputs/outputs during MAE pretraining is constrained by the limited range of color values. Therefore, reducing the magnitude of ${\bm{I}}_{\text{raw}}$ prevents exceeding these limits. However, an excessively low $r$ can result in values that are difficult to distinguish. We found that a moderate value (0.4) of $r$ performs well across most scenarios (See Section B.9 for more details). Let ${\bm{I}}_{\text{norm}}$ denote the normalized matrix, which is computed as follows:
@@
Moreover, we found that reducing the width of the visible portion can further improve performance. One possible explanation is that MAE uses a large masked ratio during pre-training, with only 25% of patches visible. Reducing the image width may align the masked ratio more closely with pre-training. Therefore, we propose multiplying $n$ by a hyperparameter $c\in[0,1]$. Similar to $r$, we found that setting $c=0.4$ performs well in our experiments (See Section B.9). This can be formulated as $n=\left\lfloor c\cdot N\cdot\nicefrac{{L}}{{(L+H)}}\right\rfloor$.
@@
Table 1: Zero-shot or few-shot results on the long-term TSF benchmark. Results are averaged across prediction lengths {96, 192, 336, 720}, with full results in Table 9 (Section B.2). Bold: the best result.
@@ 4 Experiments
Figure 4: Performance on the GIFT-Eval Leaderboard (cut-off at VisionTS’s release).
Figure 5: Aggregated results on the Monash TSF Benchmark, with full results in Table 15 (Section B.6).
We follow the standard evaluation protocol proposed by Woo et al. (2024) to test our VisionTS on 35 widely-used TSF benchmarks, and additionally evaluate it on the GIFT-Eval (Aksu et al., 2024) which is the largest TSF benchmark for zero-shot foundation models. We use MAE (Base) as our backbone by default. Baseline and benchmark details are elaborated in Section A.1.
@@ 4.1 Zero-Shot Time Series Forecasting
We first evaluate VisionTS’s zero-shot TSF performance without fine-tuning on time-series modalities. To prevent data leakage, we selected six widely-used datasets from the long-term TSF benchmark that are not included in Moirai’s pre-training set for evaluation. Since most baselines cannot perform zero-shot forecasting, we report their few-shot results by fine-tuning on the 10% of the individual target datasets. We also evaluate the Monash benchmark and GIFT-Eval benchmark. Notably, the Monash benchmark is more challenging for VisionTS since they were used in Moirai’s pre-training but not for VisionTS. We set the hyperparameters to $r=c=0.4$. Following common practice (Zhou et al., 2023; Woo et al., 2024), we conduct hyperparameter tuning on validation sets to determine the optimal context length $L$, detailed in Section B.1.
Table 2: Average MSE of different MAE variants, with full results in Table 17 (Section B.7).
@@
Table 1 shows that VisionTS surprisingly achieves the best forecasting performance in most cases (7 out of 14). Specifically, VisionTS demonstrates a relative average MSE reduction of approximately 6% compared to MoiraiSmall and MoiraiLarge, and performs comparably to MoiraiBase. When compared to the various few-shot baselines, VisionTS shows a relative average MSE reduction ranging from 8% to 84%. Given that all baselines except for VisionTS are trained on the time-series domain, this result is particularly encouraging. It suggests that the transferability from images to time-series is stronger than from text to time-series, and even comparable to the in-domain transferability between time-series. We also include a comparison with two TSF foundation models, TimesFM (Das et al., 2024) and LLMTime (Gruver et al., 2023), in Section B.3, as well as traditional algorithms (ETS, ARIMA, and Seasonal Naïve) in Section B.4. Results show that VisionTS still outperforms all of these baselines.
Results on GIFT-Eval Benchmarks
Fig. 4 shows the comparison of VisionTS with six previously published TSF foundation models on the GIFT-Eval TSF Leaderboard [^2] https://huggingface.co/spaces/Salesforce/GIFT-Eval, where VisionTS surprisingly ranked first in terms of normalized MASE. It should be noted that although some concurrent works (after the release of VisionTS) in the current leaderboard outperform VisionTS, there may be data leakage issues for these works. In contrast, visual MAE was trained on ImageNet, long before the release of GIFT-Eval leaderboard, which can ensure no leakage.
Results on Monash Benchmark
Fig. 5 shows the results aggregated from 29 Monash datasets, showing that VisionTS in the zero-shot setting surpasses all models individually trained on each dataset and significantly outperforms the other cross-domain baseline (i.e., LLMTime). It achieves second place among all baselines, just behind Moirai that pre-trained on all the training datasets. This promising result highlights VisionTS’s strong zero-shot forecasting ability and effective cross-modality transferability.
@@ 4.2 Further Analysis of VisionTS
In Table 3 (full results in Section B.7), we observe that the overall performance of three MAE variants (112M, 330M, and 657M) outperforms MoiraiSmall and MoiraiLarge. Particularly, larger models show a slight decrease in performance. This may be due to larger visual models overfitting image-specific features, reducing their transferability. A similar phenomenon was reported in Moirai, where larger models were found to degrade performance. We leave the exploration of scaling laws in image-based TSF foundation models for the future. Additionally, to explore the potential with other vision models, we also test LaMa (Suvorov et al., 2022), a visual inpainting model. Results in Section B.7 demonstrate that VisionTS with LaMa performs similarly to Moirai in the zero-shot setting. This suggests that the performance is driven by the inherent similarity between images and time series, not solely by the MAE model.
@@
Table 4: Aggregated full-shot forecasting performance on eight long-term TSF benchmarks (ETTh1, ETTh2, ETTm1, ETTm2, Illness, Weather, Traffic, and Electricity). VisionTS is fine-tuned only a single epoch on each dataset except for Illness. Due to the space limit, we report the $1^{\text{st}}$ count for each baseline, with full results in Table 21 (Section C.2).
@@
Section B.9 illustrates the impact of three hyperparameters. For context length $L$, as shown in Fig. 7, performance typically improves with increasing $L$, particularly on high-frequency datasets like Weather (10-minute frequency) and ETTm1/ETTm2 (15-minute frequency). This aligns with other TSF foundation models like Moirai. As for the normalization constant $r$ and alignment constant $c$, when both of them are around 0.4, performance is generally well across most benchmarks.
@@
We conduct experiments to validate our choices in the Alignment step, detailed in Section B.8. First, we test three different interpolation strategies, which shows that Bilinear interpolation performs best. Second, we apply horizontal and vertical flips on the image to examine whether the assumed left-to-right, top-to-bottom order is efficient. Results show that these changes do not significantly affect performance, suggesting that image reconstruction is isotropic and not influenced by certain orientation.
@@
In Appendix D, we visualize the zero-shot forecasting of VisionTS alongside the input and reconstruction images, highlighting both successful cases (where VisionTS outperforms Moirai) and failures (where Moirai prevails). When the input exhibits strong regularity (Fig. 11), VisionTS effectively forecasts both the periodicity (via segmentation) and trends (via MAE’s capabilities). In contrast, Moirai, akin to seasonal naïve methods, struggles to capture inter-period trends. For less-structured input (Figs. 12, 13 and 14), Moirai adopts a conservative approach with lower volatility to minimize errors, while VisionTS takes a more aggressive stance. This strategy occasionally yields more accurate trend predictions (Figs. 12 and 13) but may also result in greater MAE (Fig. 14).
@@ 4.3 Full-Shot Long-Term Time Series Forecasting
We evaluate the full-shot capability of each baseline trained on individual long-term TSF benchmarks. In addition to the six datasets used for zero-shot forecasting, we also include the popular Traffic and Illness datasets. As self-attention and feed-forward layers contain rich knowledge that can be transferred to TSF, we choose to fine-tune only the layer normalization (LN) layers while freezing the other parameters, which is also adopted by Zhou et al. (2023). Training details are elaborated in Section C.1.
@@
Table 4 summarizes the full-shot results, with the full results and standard deviations detailed in Section C.2. It shows that VisionTS outperforms other baselines in most cases (46 out of 80), surpassing the non-pretrained PatchTST and the language-pretrained GPT4TS. Remarkably, except for Illness with the least data, VisionTS demands only a single epoch of fine-tuning. This suggests that even minimal fine-tuning enables VisionTS to adapt to time series effectively. Compared with Table 1, fine-tuning provides limited benefits for ETTh1 and ETTh2 but significantly improves other datasets. We attribute this to the smaller data scale of ETTh1 and ETTh2.
@@
Tan et al. (2024) proposed several ablation variants for text-based foundation models, including w/o LLM (removing the LLM), LLM2Attn/LLM2Trsf (replacing the LLM with a single self-attention/Transformer layer), and RandLLM (randomly initializing the LLM). They found no significant performance differences and concluded that textual knowledge is unnecessary for TSF. We conducted similar ablations to assess the role of the vision model (VM), including w/o VM, VM2Attn, VM2Trsf, and RandVM. Section C.3 shows that these variants lead to worse performance, indicating that visual knowledge is beneficial for TSF.
@@
As stated before, we fine-tune only the layer normalization (LN). We also tested fine-tuning the bias, MLP, or attention layers, in addition to full fine-tuning and freezing. All hyperparameters were kept constant. Note that freezing differs from the previous zero-shot experiment, where a longer context length was used. Section C.3 show that fine-tuning LN is the best. Modifying MLP or attention layers results in significant performance drops, suggesting that valuable knowledge resides in these components.
@@ 5 Related Work
Self-supervised pre-training a TSF model on the same dataset used for downstream TSF tasks is a well-explored topic (Ma et al., 2023; Zhang et al., 2024), such as denoising autoencoders (Zerveas et al., 2021) or contrastive learning (Woo et al., 2022a; Yue et al., 2022). They follow a similar paradigm to the masked autoencoder (MAE) in computer vision, which is a well-studied topic in other machine learning fields, such as BERT (Devlin et al., 2019), CBraMod (Wang et al., 2025), and HuBERT (Hsu et al., 2021). However, these methods rarely examine the cross-dataset generalization capabilities. Recently, research has shifted towards training universal foundation models, by collecting large-scale time series datasets from diverse domains (Ansari et al., 2024; Goswami et al., 2024; Liu et al., 2024; Das et al., 2024; Dong et al., 2024; Feng et al., 2024) or generating numerous synthetic time series data (Fu et al., 2024; Yang et al., 2024). As a representative method, Woo et al. (2024) collected 27 billion observations across nine domains and trained TSF foundation models of various scales, achieving strong zero-shot performance. However, given the severe heterogeneity, constructing high-quality large datasets poses significant challenges for building these foundation models.